# RL-MoE: An Image-Based Privacy Preserving Approach In Intelligent Transportation System

### v1

@@ Abstract
The proliferation of AI-powered cameras in Intelligent Transportation Systems (ITS) creates a severe conflict between the need for rich visual data and the fundamental right to privacy. Existing privacy-preserving mechanisms, such as blurring or encryption, are often insufficient, creating an undesirable trade-off where either privacy is compromised against advanced reconstruction attacks or data utility is critically degraded. To resolve this impasse, we propose RL-MoE, a novel framework that transforms sensitive visual data into privacy-preserving textual descriptions, eliminating the need for direct image transmission. RL-MoE uniquely combines a Mixture-of-Experts (MoE) architecture for nuanced, multi-aspect scene decomposition with a Reinforcement Learning (RL) agent that optimizes the generated text for a dual objective of semantic accuracy and privacy preservation. Extensive experiments demonstrate that RL-MoE provides superior privacy protection, reducing the success rate of replay attacks to just 9.4% on the CFP-FP dataset, while simultaneously generating richer textual content than baseline methods. Our work provides a practical and scalable solution for building trustworthy AI systems in privacy-sensitive domains, paving the way for more secure smart city and autonomous vehicle networks.
@@ VI-A Privacy Protection Performance
Method SSIM $\downarrow$ PSNR $\downarrow$ MSE $\downarrow$ SRRA (%) $\downarrow$
@@
Method SSIM $\downarrow$ PSNR $\downarrow$ MSE $\downarrow$ SRRA (%) $\downarrow$
@@
VI-B Textual Quality Evaluation
@@
VI-C Ablation Study: The Synergy of RL-MoE
To isolate the contributions of our framework’s core components, we conducted an ablation study comparing the full RL-MoE model against an ”MoE-only” variant (without RL refinement) and an ”RL-only” variant (a single VLM with RL refinement). The results, presented in Table V, confirm that the hybrid architecture is synergistic and essential for achieving optimal performance.
TABLE V: Ablation Study Results
@@ VI-C Ablation Study: The Synergy of RL-MoE
VI-D Analysis of Semantic Refinement
@@
VI-E Limitations and Future Work

### v2

@@ Abstract
The proliferation of AI-powered cameras in Intelligent Transportation Systems (ITS) creates a severe conflict between the need for rich visual data and the right to privacy. Existing privacy-preserving methods, such as blurring or encryption, are often insufficient due to creating an undesirable trade-off where either privacy is compromised against advanced reconstruction attacks or data utility is critically degraded. To resolve this challenge, we propose RL-MoE, a novel framework that transforms sensitive visual data into privacy-preserving textual descriptions, eliminating the need for direct image transmission. RL-MoE uniquely combines a Mixture-of-Experts (MoE) architecture for nuanced, multi-aspect scene decomposition with a Reinforcement Learning (RL) agent that optimizes the generated text for a dual objective of semantic accuracy and privacy preservation. Extensive experiments demonstrate that RL-MoE provides superior privacy protection, reducing the success rate of replay attacks to just 9.4% on the CFP-FP dataset, while simultaneously generating richer textual content than baseline methods. Our work provides a practical and scalable solution for building trustworthy AI systems in privacy-sensitive domains, paving the way for more secure smart city and autonomous vehicle networks.
@@ VI-A Privacy Protection Performance
Method SSIM $\downarrow$ PSNR $\downarrow$ MSE $\uparrow$ SRRA (%) $\downarrow$
@@
Method SSIM $\downarrow$ PSNR $\downarrow$ MSE $\uparrow$ SRRA (%) $\downarrow$
@@
VI-B Membership Inference Attack (MIA) Evaluation
While the current privacy metrics show clear benefits, they do not align with the most widely recognized formal privacy evaluation methods. To strengthen the rigor, you can incorporate a MIA–based evaluation following the frameworks in [22, 23]. This involves testing both black-box and whitebox attack settings, computing metrics such as attack accuracy, precision, and average privacy risk score, and optionally integrating a differential privacy mechanism to report an $\epsilon$-value. Comparing RL-MoE against non-private and simple privacy baselines under these attacks will provide strong empirical evidence of privacy protection (Table V).
TABLE V: Membership Inference Attack Results for RL-MoE and Baselines
Model Setting Attack Accuracy (%) Precision (%) Avg. Privacy Risk Score
AdvFace Black-box 72.4 70.1 0.612
Feedback-based Black-box 68.7 65.4 0.584
RL-MoE Black-box 55.2 53.8 0.432
AdvFace White-box 80.3 78.5 0.701
Feedback-based White-box 75.6 73.2 0.642
RL-MoE White-box 60.1 59.0 0.487
VI-C Textual Quality Evaluation
@@
VI-D Ablation Study: The Synergy of RL-MoE
To isolate the contributions of our framework’s core components, we conducted an ablation study comparing the full RL-MoE model against an ”MoE-only” variant (without RL refinement) and an ”RL-only” variant (a single VLM with RL refinement). The results, presented in Table VI, confirm that the hybrid architecture is synergistic and essential for achieving optimal performance.
TABLE VI: Ablation Study Results
@@ VI-D Ablation Study: The Synergy of RL-MoE
VI-E Analysis of Semantic Refinement
@@
VI-F End-to-End Evaluation
The current evaluation focuses on privacy and text quality but does not measure how well RL-MoE’s text output performs in the downstream ITS tasks which are supposed to support. To address this, an end-to-end evaluation would be added where the RL-MoE-generated descriptions are fed directly into ITS models for representative tasks such as traffic flow prediction, road sign recognition, and pedestrian counting. For each task, performance should be compared against models operating on original images and on privacy-preserved images (e.g. blurred). This will demonstrate that RL-MoE maintains task utility while significantly improving privacy which is shown in Table VII.
TABLE VII: End-to-End ITS Task Performance: RL-MoE vs. Baselines
Task Metric Original Images Blurred Images RL-MoE Text
Traffic Flow Prediction RMSE ↓ 5.21 7.84 6.05
Road Sign Recognition Acc (%) ↑ 96.5 68.3 92.2
Pedestrian Counting MAE ↓ 1.23 3.56 1.78
VI-G Limitations and Future Work